Find the excluded values:
f(x)=3x-4/3x^2-12

Answer by ikleyn(52803)

.
Find the excluded values:
f(x)=3x-4/3x^2-12
~~~~~~~~~~~~~~~~

Excluded values are those that make the denominator equal to zero (the roots of the denominator).


In our case,  the denominator is  (3x^2 -12), and it factors this way

    3x^2 - 12 = 3*(x^2-4) = 3*(x-2)*(x+2).


So the roots are  2 and -2.  They are excluded values.      ANSWER

Solved.